OpenStudy (tygrr321):

i need help, i got half of a question and then 1 last question left plz help me i need to finish this before tomorrow

OpenStudy (tygrr321):

the equation is 4x−5(2x+1)=ax+b.

OpenStudy (tygrr321):

@jdoe0001 @pooja195

OpenStudy (jdoe0001):

thought that was done =)

OpenStudy (tygrr321):

there is still the other part of it and 1 more question

OpenStudy (jdoe0001):

hmm not sure what's expected the equation is true for x = -1/2 :)

OpenStudy (tygrr321):

and then the next part says "This means that the equation is true for (all one or no) number(s).

OpenStudy (jdoe0001):

ohhh all, one, and no well, the equation is ONLY TRUE if x = -1/2 if x = 3, no dice if x = 5, no dice if x = -9, no dice so, the equation is only true for x = -1/2 ONLY so, is true for one number/value only

OpenStudy (jdoe0001):

see it? notice, let's say, we use hmmm 7 \(4x-5(2x+1)=6x+1\qquad {\color{brown}{ x=-\frac{1}{2} }} \\ \quad \\ 4\cdot -\frac{1}{2}-10\cdot -\frac{1}{2}-5=6\cdot -\frac{1}{2}+1 \\ \quad \\ -2+5-5=-3+1\implies -2=-2\qquad \checkmark \\ \quad \\ \quad \\ 4x-5(2x+1)=6x+1\qquad {\color{brown}{ x=7 }} \\ \quad \\ 4\cdot 7-10\cdot 7+5=6\cdot 7+1 \\ \quad \\ 28-75=43\implies -47\ne 43\) notice, is true, or the equation actually EQUATES, only if x = -1/2

OpenStudy (tygrr321):

so it is one?